For O(g) at 4200 K, delta (f) G° = - 26.81 kJ/mol. For a system whose initial composition is 1.000 mol of O2(g), find the equilibrium composition at 4200 K and 3.00 bar.
1
Expert's answer
2020-11-23T07:32:59-0500

Solution.

"O2 = 2O"

"\Delta G = 2 \times (-26810) -0 = -53620 \frac{J}{mol \times K}"

"\Delta G = -RT ln(Kp)"

"Kp = exp(-\frac{\Delta G}{RT})"

Kp = 4.65

"Kp = Kn (\frac{RT}{\nu})^{\Delta \nu}"

"Kn = 1.33 \times 10^{-4}"

"1.33 \times 10^{-4} = \frac{(4x^2)}{(1-x)}"

x = 0.0057 mol

n(O) = 0.0114 mol

n(O2) = 0.9943 mol

w(O) = 1.13 %

w(O2) = 98.87 %

Answer:

w(O) = 1.13 %

w(O2) = 98.87 %